OpenStudy (calculusxy):

MEDAL! Of 20 scores, 19 are less than or equal to 10. Find the percentile rank of 10.

OpenStudy (keepcalmandqson):

what are the possible answers

OpenStudy (calculusxy):

They didn't give any options of possible answers.

OpenStudy (calculusxy):

@jhonyy9 @KamiBug @SolomonZelman @mathstudent55

OpenStudy (calculusxy):

Do I just do \(19 \div 20\) ?

OpenStudy (anonymous):

No.

OpenStudy (calculusxy):

Then?

OpenStudy (anonymous):

Give me a minute.

OpenStudy (anonymous):

Okay. A percentile rank is given by formula is: R = P/100 x (N + 1) Where. R = Percentile Rank. P = Percentile that you are trying to find. (10) N = Total number of scores Can you plug in?

OpenStudy (calculusxy):

R = 10/100 x (20 + 1) = 0.1 x 21 = 2.1

OpenStudy (anonymous):

There you go.

OpenStudy (calculusxy):

So the answer is 2.1% right?

OpenStudy (anonymous):

Yea.

OpenStudy (calculusxy):

Thank you!